Consider a C corporation. The corporation earns $7 per share before taxes. After the corporation has paid its corresponding taxes, it will distribute 100% of its earnings to its shareholders as a dividend. The corporate tax rate is 35%, the tax rate on dividend income is 28%, and the personal income tax rate is set at 28%. What are the shareholder's earnings from the corporation after all corresponding taxes are paid? Note: Express your answers in strictly numerical terms. For example, if the answer is 5%, enter 0.05 as an answer; or if the answer is $500, write enter 500 as an answer.
Answer #1

Solution:

Earning per share before taxes = $7

Corporate tax rate = 35%

Earning after tax = $7*(1-0.35) = $4.55

100% earning distribute as dividend

Tax rate on dividend income = 28%

Tax on dividend income = $4.55*(1-0.28)= $3.276

Shareholder's earning from corporation after all corresponding taxes = $3.276*(1-0.28)= $2.35872